List the first four terms of each arithmetic sequence
tn= -3+ (n-1)(5)

Answer by stanbon(75887)

List the first four terms of each arithmetic sequence
a(n) = -3+ (n-1)(5)
a(1)= -3
a(2)=-3+5=2
a(3)=-3+10=7
a(4)=-3+15=12
Cheers,
stan H.